The Coronavirus disease has led to many deaths globally. In Gauteng, the number of deaths per week increases by 9.2% every week. Suppose this trend does not change and 82 deaths have been recorded this week. Determine the number of deaths that: a) were recorded 4 weeks ago b) will be recorded in 3 weeks’ time c) will be recorded in 6 weeks’ time


1
Expert's answer
2022-02-23T14:01:37-0500

If the number of deaths this week was 82

A) Then last week "82-(9.2\%×82)=75"

2 weeks ago: "75-(9.2\%×75)=68"

3 weeks ago: "68-(9.2\%×68)=62"

4 weeks ago: "62-(9.2\%×62)=56"

The number of deaths 4 weeks ago was approximately 56 deaths.


B) if the death toll this week was 82

Next week: "82+(9.2\%×82)=90"

2 weeks time: "90+(9.2\%×90)=98"

In 3 weeks time: "98+(9.2\%×98)=107"

The expected death rate in 4 weeks time is approximately 107 deaths.


C) From the 4th week with a death toll of 107,

In 5 weeks time: "107+(9.2\%×107)=117"

In 6 weeks time: "117+(9.2\%×117)=128"

In 6 weeks time, the death rate would have increased to 128 persons